OpenStudy (anonymous):

what is 73,000,000 in scientific notaion?

OpenStudy (raffle_snaffle):



OpenStudy (raffle_snaffle):

7.3x10^7

OpenStudy (triciaal):

standard form for scientific notation a *10 ^n where 0<a < 10 and n is an integer.

OpenStudy (triciaal):

if you move the decimal to the left you are dividing so multiply by 10 to the same number of places you moved the decimal to keep the value of the number. the reverse is true.

OpenStudy (anonymous):

what is 56,4000,000,000?

OpenStudy (triciaal):

Given 73,000,000 (the decimal point is at the end of the whole number) to get the number between one and ten you moved the decimal seven places to the left 10^-7 to keep the same value of the number multiply by 10^7. 10^-7* 10^7 = 10^-7+7 = 10^0 any number raised to the zero = 1 so 73,000,000 = 7.3*10^7

OpenStudy (anonymous):

what?

OpenStudy (triciaal):

follow what I did I'll be here

OpenStudy (triciaal):

ok let's do it together slowly

OpenStudy (triciaal):

56,4000,000,000 where is the decimal point?

OpenStudy (anonymous):

there is none

OpenStudy (triciaal):

read my post above slowly

OpenStudy (triciaal):

example 6 = 6.0

OpenStudy (triciaal):

which way do you need to move the point so that the first digit is less than 10 and more than 0? how many places?

OpenStudy (triciaal):

are you still reading or do you have questions? I can only help you if you talk to me. I am here to make sure you can do the problem

OpenStudy (anonymous):

what is 564,000,000,000 in scientific notaion?

OpenStudy (triciaal):

(the decimal point is at the end of the whole number) to get the number between one and ten you moved the decimal

OpenStudy (triciaal):

how many places do you need to move the decimal and what direction?

OpenStudy (triciaal):

do you still need help?

OpenStudy (anonymous):

no